# A Comparison of Precinct and District Voting Data Using Persistent Homology to Identify Gerrymandering in North Carolina

### v3

@@ Abstract
We present an extension of Feng & Porter’s 2019 paper on the use of the level-set method for the construction of a filtered simplicial complex from geospatial election data. Precincts are regarded to be too small to be gerrymandered, allowing us to identify discrepancies between precinct and district level voting data to quantify gerrymandering in the United States.
Comparing the persistent homologies of democratic voting areas on the precinct and district level shows when areas have been ‘cracked’ or ‘packed’ for partisan gain.
This analysis was done for North Carolina House of Representatives elections (2012-2024). NC has been redistricted 4 times in the past 10 years, whereas most states redistrict decennially, allowing us to understand how and when redistricted maps deviate from precinct-level voting data, and when gerrymandering occurs. Comparing persistence barcodes at the precinct and district levels (using the bottleneck distance) shows that precinct-level voting patterns do not significantly fluctuate biannually, while district level patterns do, suggesting that shifts are likely a result of redistricting rather than voter behavior, providing strong evidence of gerrymandering.
NC Election data was collected from the public domain. Composite shapefiles were created using QGIS and R, and rasterized using Python. The level-set method was employed to generate filtered similar complexes. Persistence barcodes were produced using GUDHI and PHAT libraries.
Additionally, we compare our results with traditional measures such as Polsby-Popper and Reock scores (gerrymandering identification measures). This research presents a novel application of topological data analysis in evaluating gerrymandering.
 [^1] This work was presented at the 39th Summer Topology and Its Applications Conference – Topology and Computing (August 11, University of South Alabama); will be presented at the 57th Annual Northeastern Political Science Association meeting – American Politics (November 6-8, 2025, Philadelphia, PA); and at the Social Choice: Theory and Computation Conference – Poster Session (October 15-17, 2025, Wellesley College).

### v4

@@ Abstract
Gerrymandering is one of the biggest threats to American democracy. By manipulating district lines, politicians effectively choose their voters, rather than the other way around. Current gerrymandering identification methods (namely the Polsby-Poppers and Reock scores) focus on the compactness of congressional districts, making them extremely sensitive to physical geography.
To address this gap, we extend Feng & Porter’s 2021 paper, which used the level-set method to turn geographic shapefiles into filtered simplicial complexes, in order to compare precinct level voting data to district level voting data. As precincts are regarded as too small to be gerrymandered, we are able to identify discrepancies between precinct and district level voting data to quantify gerrymandering in the United States.
By comparing the persistent homologies of Democratic voting regions at the precinct and district levels, we detect when areas have been ‘cracked’ (split across multiple districts) or ‘packed’ (compressed into one district) for partisan gain.
This analysis was conducted for North Carolina House of Representatives elections (2012-2024). NC has been redistricted 4 times in the past 10 years, unusually frequent as most states redistrict decennially, making it a valuable case study.
By comparing persistence barcodes at the precinct and district levels (using the bottleneck distance), shows that precinct-level voting patterns do not significantly fluctuate biannually, while district level patterns do, suggesting that shifts are likely a result of redistricting rather than voter behavior, providing strong evidence of gerrymandering.
This research presents a novel application of topological data analysis in evaluating gerrymandering, and shows persistent homology can be useful in discerning gerrymandered districts.
@@ 8  Acknowledgments
This work is supported by the Carnegie Foundation and Carnegie Young Leaders. Thank you for your support.